# Minimizing Selection Bias in Inclusive Jets in Heavy-Ion Collisions with Energy Correlators — arXiv note Carlota Andres

### v2

@@ Abstract
This note serves as a companion to a Letter Andres et al. (2024a), where we introduce a new energy correlator-based observable designed to minimize the impact of selection bias due to energy loss in inclusive jets in heavy-ion collisions. Here, we apply the method outlined in the Letter to the first-ever measurement of energy correlators in heavy-ion collisions, recently released by the CMS Collaboration Viinikainen (2024); CMS (2024).
@@
This brief note applies the method introduced in Andres et al. (2024a) to the CMS measurement of the two-point energy correlator (E2C) in inclusive jets in heavy-ion (A-A) collisions Viinikainen (2024); CMS (2024). In Andres et al. (2024a), we introduced a new correlator-based observable, defined as
where denotes the E2C distribution Chen et al. (2020); Komiske et al. (2023) in heavy-ion jets, with being the angle between detectors, and is the so-called unbiasing function, designed to correct for leading-order energy loss effects the E2C distribution. For the derivation and explicit functional form of the unbiasing function, we refer the reader to Andres et al. (2024a). Once the E2C distribution is measured, the unbiasing function can be directly obtained. We note that in proton-proton (p-p) collisions, is unity by construction, making identical to E2C for p-p jets.
@@
In Fig. 1, we present a comparison of the E2C (left panel) and E2C/ (right panel) for inclusive jets with reconstructed in the range GeV. Results are shown for both p-p and the 0-10 centrality class of Pb-Pb collisions. In the top left panel, we observe that the Pb-Pb E2C (green) is shifted toward smaller values compared to the p-p E2C (black). This shift results in the overall suppression of the Pb-Pb spectrum relative to the p-p one for , as illustrated in the bottom left panel. When the unbiasing function is applied (right panel), the suppression due to energy loss is significantly reduced. Moreover, the ratio between the Pb-Pb E2C/ and the p-p one (right bottom panel) becomes completely flat in the hadronization region (). This shows that, within the current uncertainties, the observed modifications in the E2C in Pb-Pb compared to p-p in the deep hadronization region (left panel, ) are fully attributable to selection bias effects. In the large angular region (), we observe that the enhancement with respect to p-p is larger for the E2C/ distribution than for the E2C, indicating that the application of the unbiasing function has effectively mitigated energy loss effects in this region. However, a quantitative description of the remaining enhancement in the E2C/ spectrum requires further investigation, as several phenomena—such as medium-induced radiation and medium response —are predicted to contribute to this behavior Andres et al. (2023a, b, 2024b); Bossi et al. (2024); Yang et al. (2024).
@@
In Fig. 2, we analyze the centrality dependence of the E2C and E2C/ spectra. The filled squares (open circles) represent the E2C (E2C/) within inclusive jets with reconstructed in the range GeV divided by the corresponding p-p baseline. Each panel corresponds to a different centrality class in Pb-Pb collisions. It is evident that as centrality decreases, the amplitude of the wide-angle enhancement in both the E2C and E2C/ ratios diminishes. This observation aligns with the general prediction that both selection bias due to energy loss and other medium effects are reduced in more peripheral collisions Andres et al. (2023a, b, 2024b); Bossi et al. (2024); Viinikainen (2024). The - centrality class is particularly noteworthy. Before applying the unbiasing function (filled dark blue squares), the Pb-Pb/p-p ratio within this class appears to display a nontrivial structure. While, after applying the unbiasing function, the E2C/ Pb-Pb/p-p ratio (empty light blue circles) becomes completely flat. This indicates that a small selection bias was the sole cause of deviation from the vacuum baseline and that the unbiasing function has accurately corrected for this effect.
@@
Fig. 3 focuses on the -dependence of the results. The filled squares (empty circles) represent the E2C (E2C/) distribution within inclusive jets in the 0-10 centrality class in Pb-Pb collisions divided by the E2C distribution in inclusive p-p jets. Each panel corresponds to a different range of reconstructed jet transverse momenta. It can be clearly observed that the wide-angle enhancement decreases in amplitude as the jet increases. Additionally, the wide-angle structure shifts to slightly smaller values as the jet increases. These qualitative behaviors were predicted as consequences of medium-induced radiation prior to the unveiling of the CMS measurement Andres et al. (2023a, b, 2024b). However, a comprehensive analysis, including the extraction of scaling laws governing this enhancement, is required to disentangle medium-induced radiation from other effects, such as medium response Yang et al. (2024); Bossi et al. (2024), and to identify possible signatures of color coherence dynamics Andres et al. (2023a, b).
@@
Finally, we present in Fig. 4 the evaluation of the following approximate formula, derived in Andres et al. (2024a),
using the CMS p-p data on the two-point energy correlator within p-p jets CMS (2024) and for . We refer the interested reader to Andres et al. (2024a) for further details on this equation and the role it plays in the derivation of the unbiasing function . The derivative on the right hand side of eq. (2) was evaluated analytically, while that on the left hand side was computed numerically. We did not compute error bars for the numerical derivatives, as the bin-to-bin correlated uncertainties of the CMS measurement are not yet available. The results shown in Fig. 4 show good agreement between both sides of eq. (2) over the complete angular range studied, even around the hadronization transition at .
@@
Conclusions—This companion note to the Letter Andres et al. (2024a) presents results on the correlator-based observable introduced in Andres et al. (2024a), E2C/, using the CMS measurement of the two-point energy (E2C) correlator within inclusive jets in Pb-Pb collisions at the LHC Viinikainen (2024); CMS (2024). The E2C/ observable shows significantly reduced sensitivity to energy loss effects compared to the corresponding E2C for jets with transverse momenta between 120 GeV and 200 GeV, across different centrality classes in Pb-Pb collisions. Our current analysis does not account for correlations in the error propagation, and thus we encourage the experiments to perform this measurement.

### v3

@@ Abstract
This note serves as a companion to a Letter Andres et al. (2025), where we introduce a new energy correlator-based observable designed to minimize the impact of selection bias due to energy loss in inclusive jets in heavy-ion collisions. Here, we apply the method outlined in the Letter to the first-ever measurement of energy correlators in heavy-ion collisions, recently released by the CMS Collaboration Viinikainen (2024); CMS (2024).
@@
This brief note applies the method introduced in Andres et al. (2025) to the CMS measurement of the two-point energy correlator (E2C) in inclusive jets in heavy-ion (A-A) collisions Viinikainen (2024); CMS (2024). In Andres et al. (2025), we introduced a new correlator-based observable, defined as
where denotes the E2C distribution Chen et al. (2020); Komiske et al. (2023) in heavy-ion jets, with being the angle between detectors, and is the so-called unbiasing function, designed to correct for leading-order energy loss effects the E2C distribution. For the derivation and explicit functional form of the unbiasing function, we refer the reader to Andres et al. (2025). Once the E2C distribution is measured, the unbiasing function can be directly obtained. We note that in proton-proton (p-p) collisions, is unity by construction, making identical to E2C for p-p jets.
@@
In Fig. 1, we present a comparison of the E2C (left panel) and E2C/ (right panel) for inclusive jets with reconstructed in the range GeV. Results are shown for both p-p and the 0-10 centrality class of Pb-Pb collisions. In the top left panel, we observe that the Pb-Pb E2C (green) is shifted toward smaller values compared to the p-p E2C (black). This shift results in the overall suppression of the Pb-Pb spectrum relative to the p-p one for , as illustrated in the bottom left panel. When the unbiasing function is applied (right panel), the suppression due to energy loss is significantly reduced. Moreover, the ratio between the Pb-Pb E2C/ and the p-p one (right bottom panel) becomes completely flat in the hadronization region (). This shows that, within the current uncertainties, the observed modifications in the E2C in Pb-Pb compared to p-p in the deep hadronization region (left panel, ) are fully attributable to selection bias effects. In the large angular region (), we observe that the enhancement with respect to p-p is larger for the E2C/ distribution than for the E2C, indicating that the application of the unbiasing function has effectively mitigated energy loss effects in this region. However, a quantitative description of the remaining enhancement in the E2C/ spectrum requires further investigation, as several phenomena—such as medium-induced radiation and medium response —are predicted to contribute to this behavior Andres et al. (2023a, b, 2024); Bossi et al. (2024); Yang et al. (2024).
@@
In Fig. 2, we analyze the centrality dependence of the E2C and E2C/ spectra. The filled squares (open circles) represent the E2C (E2C/) within inclusive jets with reconstructed in the range GeV divided by the corresponding p-p baseline. Each panel corresponds to a different centrality class in Pb-Pb collisions. It is evident that as centrality decreases, the amplitude of the wide-angle enhancement in both the E2C and E2C/ ratios diminishes. This observation aligns with the general prediction that both selection bias due to energy loss and other medium effects are reduced in more peripheral collisions Andres et al. (2023a, b, 2024); Bossi et al. (2024); Viinikainen (2024). The - centrality class is particularly noteworthy. Before applying the unbiasing function (filled dark blue squares), the Pb-Pb/p-p ratio within this class appears to display a nontrivial structure. While, after applying the unbiasing function, the E2C/ Pb-Pb/p-p ratio (empty light blue circles) becomes completely flat. This indicates that a small selection bias was the sole cause of deviation from the vacuum baseline and that the unbiasing function has accurately corrected for this effect.
@@
Fig. 3 focuses on the -dependence of the results. The filled squares (empty circles) represent the E2C (E2C/) distribution within inclusive jets in the 0-10 centrality class in Pb-Pb collisions divided by the E2C distribution in inclusive p-p jets. Each panel corresponds to a different range of reconstructed jet transverse momenta. It can be clearly observed that the wide-angle enhancement decreases in amplitude as the jet increases. Additionally, the wide-angle structure shifts to slightly smaller values as the jet increases. These qualitative behaviors were predicted as consequences of medium-induced radiation prior to the unveiling of the CMS measurement Andres et al. (2023a, b, 2024). However, a comprehensive analysis, including the extraction of scaling laws governing this enhancement, is required to disentangle medium-induced radiation from other effects, such as medium response Yang et al. (2024); Bossi et al. (2024), and to identify possible signatures of color coherence dynamics Andres et al. (2023a, b).
@@
Finally, we present in Fig. 4 the evaluation of the following approximate formula, derived in Andres et al. (2025),
using the CMS p-p data on the two-point energy correlator within p-p jets CMS (2024) and for . We refer the interested reader to Andres et al. (2025) for further details on this equation and the role it plays in the derivation of the unbiasing function . The derivative on the right hand side of eq. (2) was evaluated analytically, while that on the left hand side was computed numerically. The error bands on the numerical derivatives account for uncertainties from both the numerical method used for their computation and the experimental data. Experimental uncertainties are treated as uncorrelated, as bin-to-bin correlations from the CMS measurement are not yet available. Notably, the overall uncertainty is dominated by the numerical method. The results shown in Fig. 4 show good agreement between both sides of eq. (2) over the complete angular range studied, even around the hadronization transition at .
@@
Conclusions—This companion note to the Letter Andres et al. (2025) presents results on the correlator-based observable introduced in Andres et al. (2025), E2C/, using the CMS measurement of the two-point energy (E2C) correlator within inclusive jets in Pb-Pb collisions at the LHC Viinikainen (2024); CMS (2024). The E2C/ observable shows significantly reduced sensitivity to energy loss effects compared to the corresponding E2C for jets with transverse momenta between 120 GeV and 200 GeV, across different centrality classes in Pb-Pb collisions. Our current analysis does not account for correlations in the error propagation, and thus we encourage the experiments to perform this measurement.